the density of DNA is 1.1gram per mL and molecular weight is 6000gram per mole.what is the value of average volume covered by its single molecule?
1
Expert's answer
2017-06-08T11:57:12-0400
d(DNA)=1.1 g/mL
M(DNA)=6000 g/mol

If we divide molecular weight of the DNA molecule by density we will get the volume of 1 mole of DNA:
V=(M(DNA))/(D(DNA))=(6000 G/MOL)/(1.1 G/ML)=5454 ML/MOL